OpenStudy (anonymous):

Need help writing an equation: A TANK IS FILLING WITH WATER FROM A NATURAL SPRING. TWO DAYS AGO THE WATER WAS 10 FEET DEPTH, AND YESTERDAY THE WATER WAS 12 FEET. ASSUME THAT THE WATER DEPTH CONTINUES TO RISE AT THIS RATE AFTER TODAY.

OpenStudy (anonymous):

the slope is 2 thats what I already know and im sure :)

OpenStudy (anonymous):

x+2 let x be the day

OpenStudy (anonymous):

am i rite!!!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

which is y=mx+b right?

OpenStudy (anonymous):

what coordinates do you use?

OpenStudy (anonymous):

so -2=2(2)+b is the 2 in parenthesis be a negative or positive

OpenStudy (anonymous):

what am I suppose to do??????????????

OpenStudy (anonymous):

No

OpenStudy (anonymous):

whats a whats d

OpenStudy (anonymous):

help me im confused

OpenStudy (anonymous):

I SERIUOSLY NEED HELP CAN YOU JUST SHOW MW HOW TO FIND THE ANSWER! I LEARN FROM THE WORK

OpenStudy (anonymous):

SO YOU DISTRIBUTE RIGHT

OpenStudy (anonymous):

so the equation is y=2n+12

OpenStudy (anonymous):

I keep putting that in the answer and it says Im wrong

OpenStudy (dumbcow):

hmm i think it would be y = 2x+14 since Today the water level is 14

OpenStudy (anonymous):

Let me try that kay?

OpenStudy (anonymous):

nope its wrong

OpenStudy (dumbcow):

or using point slope form: point (-1,12) slope = 2 y-12 = 2(x-(-1)) y -12 = 2x +2 y = 2x +14

OpenStudy (anonymous):

l=2d+10. l= water level and d= days

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Want an explanation?

OpenStudy (anonymous):

NO THANKS

OpenStudy (anonymous):

k, once someone ahelps you with the problem the answer will be l=2d+10. l= water level and d= days ;)

OpenStudy (dumbcow):

10,12,14 ... the rate is increase of 2 per day so the equation is: y = 2x +b to know "b" we need what day is the starting point i was assuming x=0 meant TODAY is there more info in the problem ?

OpenStudy (anonymous):

sorry wait tO FIND b IN Y=2X+B YOU CAN SUBSTITUTE IN ONE OF THE COORDINATES PAIRS -2 for x and 10for y. for example you can substitute -2 for x and 10 for y. this gives 10=(2)(-2)+b. solve this equation for b and substitue that value for in to the expression

OpenStudy (anonymous):

that what that says

OpenStudy (mathmate):

I didn't follow all the conversations. Now that we should be able to make a rule, what is the question? I.e. what are we required to find?

OpenStudy (mathmate):

We need the water level today?

OpenStudy (mathmate):

or the rule?

OpenStudy (anonymous):

an equation

OpenStudy (anonymous):

ill atach the question so you can see it okay?

OpenStudy (mathmate):

Like dc mentioned, it depends on where we take x=0. The rule (equation) is different if we take x=0 as today, yesterday or the day before. So there are three possibilities. How many have you already tried?

OpenStudy (anonymous):

this is the question

OpenStudy (mathmate):

It is much simpler if you post the question, because in the question it says "Time elapsed from today" which means that today is the same as x=0. The complete question is: A TANK IS FILLING WITH WATER FROM A NATURAL SPRING. TWO DAYS AGO THE WATER WAS 10 FEET DEPTH, AND YESTERDAY THE WATER WAS 12 FEET. ASSUME THAT THE WATER DEPTH CONTINUES TO RISE AT THIS RATE AFTER TODAY. 1. When will the water be 14 feet deep? 2. How long will the water tale to reach a depth of 50 feet? 3. How deep will the water in the tank be in 21 days? 4. How deep was the water in the tank three days ago? Time elapsed from today The depth of the water in the tank Unit: days Unit: feet Expression: x Expression: IDK Question 1: Question 1: 14 Question2: Question2: 50 Question3: 21 Question3: Question 4: -3 Question 4: We will answer them one by one.

OpenStudy (anonymous):

THE EQAUTION IS JUST 2X+14 NO Y= 'S

OpenStudy (mathmate):

If you make a table, x=days from today, H=height x H -2 10 -1 12 So the rate of increase is 2 feet per day, so today, the height is 14. The rule is therefore H=2x+14 (today is day zero) Is that ok so far?

OpenStudy (mathmate):

An equation always needs an equal sign. If you say the rule, it could be 2x+14 (with height being understood).

OpenStudy (mathmate):

For questions 2, 3 and 4, you need to substitute the known values of x (or H) into the equation, and solve for the other variable. For example, if it asks "when will the water level be 26 feet high", then put H=26 into the above equation ans solve for x (days from today): 26=2x+14 => 26-14=2x+14-14 => 2x=12 => x=6 So 6 days from now, the water level will be 26 feet high. etc.

OpenStudy (anonymous):

I got it thanks

OpenStudy (mathmate):

yw! :)

OpenStudy (anonymous):

Gave you a medal for your hard work mathmate.

OpenStudy (mathmate):

Thank's a million! :)